### v1

# AgentArch: A Comprehensive Benchmark to Evaluate Agent Architectures in Enterprise

Abstract
While individual components of agentic architectures have been studied in isolation, there remains limited empirical understanding of how different design dimensions interact within complex multi-agent systems. This study aims to address these gaps by providing a comprehensive enterprise-specific benchmark evaluating 18 distinct agentic configurations across state-of-the-art large language models. We examine four critical agentic system dimensions: orchestration strategy, agent prompt implementation (ReAct versus function calling), memory architecture, and thinking tool integration. Our benchmark reveals significant model-specific architectural preferences that challenge the prevalent one-size-fits-all paradigm in agentic AI systems. It also reveals significant weaknesses in overall agentic performance on enterprise tasks with the highest scoring models achieving a maximum of only 35.3% success on the more complex task and 70.8% on the simpler task. We hope these findings inform the design of future agentic systems by enabling more empirically backed decisions regarding architectural components and model selection.  [^1] GitHub Repository: https://github.com/ServiceNow/AgentArch
1 Introduction
Large language models (LLMs) have evolved beyond simple text generation to function as autonomous agents capable of iterative decision-making and complex task completion (Yao et al. 2025). This transformation has opened new possibilities for enterprise automation, where agentic AI systems can interact with existing workflows, databases, and business processes to accomplish multi-step objectives with minimal human intervention.
Despite growing interest in agentic systems, practitioners face two critical gaps: existing research predominantly evaluates individual components in isolation—orchestration strategies, prompting approaches, memory management, etc—without examining how these architectural choices interact with each other in real-world scenarios. Additionally, while existing benchmarks span diverse domains from gaming to research tasks, few focus specifically on enterprise workflows. Enterprise use cases are uniquely constrained by the need for reliability and seamless integration with existing business systems. The fragmented approach of prior benchmarks leaves enterprise developers without empirical guidance for selecting optimal combinations of architectural components or choosing appropriate models for their specific set ups.
This paper presents a comprehensive evaluation of 18 distinct agentic architectures across state-of-the-art LLMs, systematically examining the interactions between four key architectural dimensions: orchestration strategy (single-agent vs. multi-agent), agent style (function calling vs. ReAct), memory management (complete vs. summary), and thinking tool integration (Yao et al. 2023). We test each architectural combination on two enterprise use cases of varying complexity, providing practical insights for agentic system designers.
Our findings reveal significant model-specific architectural preferences, challenging assumptions about universally optimal designs. We find that even between use cases, models perform best on different architectures. While larger models display more robust performance across architectures, we find that on simpler tasks, smaller models can match performance under their best performing architecture. We also observe significant weaknesses across models when ReAct prompting is used in multi-agent systems. Our results provide empirically-grounded guidance for practitioners building enterprise agentic systems, enabling more informed decisions about architectural components and model selection based on specific use case requirements.
2 Related Work
Requesting Time Off (TO) Customer Request Routing (CR)
Description Simple, structured workflow for PTO eligibility verification and request processing. Tests basic multi-step reasoning with clear success criteria. Complex workflow mimicking intelligent customer service systems. Handles simple requests automatically while escalating complex issues. Tests classification and escalation decisions.
Tools 8 custom tools 31 custom tools
Agents 3 agents 9 agents
Agent Types Employee Information Retriever Agent, Balance Eligibility Agent, Leave Approval Agent Request Validation Agent, Duplicate Case Detection Agent, Case Creation Agent, Informational Queries Agent, Transactional Queries Agent, Intent and Sentiment Extraction Agent, Entity Extraction Agent, Document Verification Agent, Email Generation Agent
Key Challenges Date calculations; leave balance verification; policy compliance; multi-step approval workflow Appropriate escalation decisions; context preservation; handling ambiguous requests; complex routing logic; instruction following
Table 1: Detailed comparison of the two enterprise use cases evaluated in our benchmark.
The evaluation of agentic AI systems has emerged as a critical research area, with existing benchmarks examining various components of agent architectures. However, these efforts have predominantly focused on isolated aspects of agentic systems rather than their integrated performance.
Tool calling is perhaps the most fundamental aspect of all agentic systems. Existing tool calling benchmarks (Liu et al. 2024b; Patil et al. 2025; Wu et al. 2024) have revealed significant LLM weaknesses in key aspects of tool calling including tool selection accuracy, tool input accuracy, valid output formatting, and issues with latency and unnecessary repetition of tool calls. Additional benchmarks focus on more complex scenarios finding that LLM performance on tool calling decreases when combined with complex instruction following (Yao et al. 2025) and multi step reasoning in nested tool calling scenarios (Basu et al. 2025).
Another key aspect of agentic systems is memory and context management. LLMs are inherently limited by their context windows which provides a unique challenge for those building agentic systems to ensure that the LLMs keep the most relevant information in context. This is relevant for both conversational and tool-use settings, as conversational context accumulates over time, while tool calls can significantly increase context length due to their frequency or output size (Maharana et al. 2024a). Existing research presents a variety of mechanisms for improving memory management in agentic systems such as treating agent memory as a self directed operating system with different storage tiers (Packer et al. 2024) , creating dynamic knowledge networks for connecting information (Xu et al. 2025), and implementing a dual-track memory system meant to resemble long-term and short-term memory in the human brain (Liu et al. 2024a).
As LLMs enable more complex agentic workflows, Multi-Agent Systems (MAS) have emerged to support use cases involving specialized agents, prompting the development of benchmarks to evaluate their effectiveness. (Zhu et al. 2025) examine LLM performance across various MAS structures, including hierarchical and distributed planning, evaluating task completion and planning capabilities to assess LLM effectiveness in both planning and execution roles. (Agashe, Fan, and Wang 2025) and (Geng and Chang 2025) focus specifically on evaluating planning and coordination between agents in MAS finding that planning abilities are a key component of success in MAS. (Yue et al. 2025) illustrate how multi-agent systems can benefit from routing frameworks that help ensure that the appropriate agents, communication strategies, and LLMs are selected for a given task. Similarly, (Hou, Tang, and Wang 2025) introduce a dynamic MAS framework that incorporates hierarchical planning along with a search process for selecting the optimal multi-agent configuration for the task at hand. Other work on dynamically generated MAS demonstrates how scaling single agents into MAS can lead to performance increases (Yuan et al. 2025). Recent work compares heterogeneous LLMs versus a single model in MAS, highlighting how MAS can be used to leverage individual model strengths (Ye et al. 2025).
As our work is primarily focused on enterprise benchmarks, we also explore existing work on agents in enterprise domains. Enterprise-focused evaluation has consistently revealed a gap between existing academic benchmarks and real-world applicability, with even state-of-the-art models struggling on simple workplace tasks (Styles et al. 2024). Benchmarks across different enterprise domains consistently show that realistic business task complexity significantly exceeds what current models can handle reliably, with performance typically degrading in multi-turn interactions and interconnected tasks (Huang et al. 2025a, b; Drouin et al. 2024; Styles et al. 2024).
While prior benchmarks have advanced understanding of individual agentic components, they remain fragmented—evaluating tool use (Yao et al. 2025; Patil et al. 2025) (Basu et al. 2025), memory (Maharana et al. 2024b), and coordination (Geng and Chang 2025; Agashe, Fan, and Wang 2025) in isolation. Existing enterprise benchmarks address task complexity but lack systematic analysis of how architectural choices interact across scenarios. As a result, developers lack empirical guidance for selecting and combining agentic components under real-world enterprise constraints.
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 36.9 21.2 6.2 0.0 7.1 0.0 1.2 0.0 0.4 0.0 28.1 0.0
Complete Disabled 31.9 17.5 4.8 1.2 7.3 0.5 1.2 0.0 0.0 0.0 28.7 0.2
Function Calling Complete Enabled 58.8 14.8 50.2 4.2 65.8 4.8 4.6 20.5 0.0 0.0 67.7 15.2
Complete Disabled 54.4 16.2 46.5 5.0 56.0 4.8 5.4 21.8 0.0 0.0 57.7 17.2
Summarized Enabled 51.9 15.2 49.0 0.8 59.0 2.5 5.0 18.5 0.0 0.0 59.6 16.8
Summarized Disabled 52.7 15.8 45.0 0.8 49.6 1.0 4.8 19.0 0.0 0.0 44.2 16.5
Orch-Isolated React Complete Enabled 28.7 17.2 3.8 0.0 4.6 0.0 2.3 0.0 0.2 0.0 28.3 0.0
Complete Disabled 25.8 16.0 5.8 0.2 5.6 0.0 1.5 0.0 0.0 0.0 28.1 0.5
Function Calling Complete Enabled 58.8 17.2 53.5 1.8 67.1 0.8 6.0 22.2 0.0 0.0 68.5 14.8
Complete Disabled 54.6 16.2 52.1 1.8 57.1 1.5 4.0 21.0 0.0 0.0 57.5 16.2
Summarized Enabled 54.6 16.5 46.0 1.8 59.6 2.0 2.9 20.5 0.0 0.0 58.1 16.8
Summarized Disabled 51.5 15.8 44.2 3.0 54.4 1.5 4.8 21.2 0.0 0.0 44.8 14.5
Single Agent React Complete Enabled 34.4 1.0 15.8 0.2 51.2 1.2 8.8 1.5 6.2 0.0 63.7 3.5
Complete Disabled 36.7 3.8 9.6 0.0 41.5 1.2 2.9 2.2 12.3 0.0 52.7 9.0
Function Calling Complete Enabled 68.1 22.2 39.0 3.0 35.4 0.2 56.7 1.5 0.0 0.0 66.7 33.8
Complete Disabled 49.6 21.2 26.9 3.5 19.8 0.2 55.8 2.0 3.8 0.0 31.7 35.0
Summarized Enabled 70.8 21.5 37.9 2.5 33.5 0.5 55.0 1.5 0.6 0.0 64.4 33.2
Summarized Disabled 48.5 19.8 28.7 2.8 24.4 0.5 55.8 1.0 0.0 0.0 31.7 35.2
Table 2: Pass@1 Acceptable Rate. Bolded scores indicate highest score per model on use case.
Model TO CR
Mean SD CV Mean SD CV
GPT-4.1 48.2 13.0 27.0 16.1 5.5 34.4
GPT-4o 31.4 18.7 59.7 1.8 1.5 83.9
GPT-4.1-mini 38.8 22.2 57.2 1.3 1.5 110.7
o3-mini 15.5 22.3 143.7 9.7 10.1 104.0
LLaMA 70B 1.1 3.1 286.8 0.0 0.0 0.0
Sonnet 4 49.0 15.7 32.1 15.5 12.3 79.1
Table 3: Model consistency analysis - mean, standard deviation, and coefficient of variance for acceptable pass@1 scores. Performance statistics across all 18 architectural configurations for each use case.
3 Methodology
3.1 Experiment Set Up
Use Cases We evaluate two enterprise workflows of varying complexity to examine how task complexity affects agentic performance on real-world enterprise automation scenarios. Both of these use cases involve specific directions to complete tasks in a specified order, reflecting the rigid nature of many enterprise workflows where steps must take place in a predefined order. For each request in a use case, we define an expected final decision (e.g. leave approve or rejected). We create mock data so that each tool request returns a deterministic result. Unlike existing benchmarks that use simplified, clean responses, we deliberately construct enterprise-realistic data that is complex, lengthy, and messy. Knowledge base articles span thousands of words with extensive documentation reflecting real enterprise systems. Tool responses return complex JSON objects where relevant information is often buried within metadata, error codes, and auxiliary fields. This approach better represents the information processing challenges agents face in production environments where data is rarely clean or perfectly structured.
Figure 1: Orchestration, Isolated Agents. Agents ask orchestrator for help and orchestrator selects agent to assist.
Figure 2: Orchestration, Open Agent Network. Agents ask other agents for help directly.
Requesting Time Off (TO) represents a simple workflow involving PTO eligibility verification and request processing. This use case tests basic multi-step reasoning, counting, and decision-making with clear success criteria, employing 8 custom tools and 3 agents.
Customer Request Routing (CR) represents a complex workflow that mimics intelligent customer service systems, handling simple requests automatically while escalating complex issues to human agents. This tests nuanced classification, context understanding, and appropriate escalation decisions with 31 custom tools and 9 agents. Each use case includes 60 user utterances designed to cover the full spectrum of real-world requests, including successful scenarios, edge cases, and failure conditions that would occur in actual enterprise deployments. Implementation details are provided in Appendix A.1.
Orchestration Strategies. We evaluate four orchestration strategies to examine single-agent versus multi-agent performance across key dimensions: decision-making with varying option sets, collaboration costs and benefits, and specialist versus generalist agent effectiveness. 1) Orchestrator-led, Isolated Agent. An orchestrator assigns tasks to specialist agents and controls all inter-agent communication. When agents need additional information, they request help through the orchestrator, who selects an appropriate agent to respond. This tests whether centralized planning and controlled information flow improve outcomes. 2) Orchestrator-led, Open Agent Network. The orchestrator assigns initial tasks, but agents communicate directly with each other when additional information is needed. Agents select collaborators independently using agent descriptions. This examines whether direct agent-to-agent communication is more efficient than orchestrator-mediated interaction. 3) Single Agent. A single agent receives the user request and has access to all available tools, with no collaboration mechanisms. This tests agent performance in large, unfiltered option spaces.
Agents. We evaluate two prevalent agent styles to examine how different prompting paradigms influence model performance. With function calling, the model directly selects tools from the available toolkit, leveraging inherent function calling capabilities without explicit intermediate reasoning. Using ReAct, The model is instructed to output its thoughts before selecting actions thus using a structured reasoning-action framework (Yao et al. 2023). We investigate whether the explicit reasoning in ReAct prompting yields substantially different outcomes compared to direct function calling. Implementation details and prompts are provided in the appendix.
Memory Management. We evaluate two information sharing approaches to examine the trade-off between context completeness and input context length. With complete memory, agents see all previous tool calls, parameters, and responses from other agents, maximizing context but increasing prompt length. With summarized memory, Agents receive only final summaries from previous agents, reducing context length while potentially limiting coordination effectiveness. This tests whether comprehensive information sharing improves performance or whether condensed summaries provide sufficient context for effective collaboration.
Thinking Tools. We evaluate whether providing explicit reasoning tools improves model performance on multi-step analysis tasks. When thinking tools are enabled, models receive access to math for calculations and synthesize_collected_information for data organization, particularly relevant for PTO eligibility where models must calculate date ranges and compare leave balances. We hypothesize that models without built-in reasoning capabilities will benefit from these structured tools. See appendix A.2 for more details and examples.
Model Settings. We evaluate our benchmark using six models: OpenAI GPT-4.1, OpenAI-GPT-4o, LLaMA 3.3 70B, OpenAI GPT-4.1-mini, Anthropic Claude Sonnet 4, and OpenAI o3-mini. We evaluated Claude Sonnet 4 with reasoning disabled. For all experiments, we set temperature to 0 and use no other sampling parameters. We selected this set of models to represent state-of-the-art models of varying sizes, including models with and without in-built reasoning capabilities. For function calling architectures, we use the in-built OpenAI function calling APIs for OpenAI models and LLaMa 3.3. For Sonnet 4, we use the Bedrock API.
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 1.0 20.8 19.4 77.2 15.2 64.8 64.0 96.5 29.2 35.6 0.0 31.0
Complete Disabled 2.1 18.8 22.5 71.5 18.3 65.2 67.5 98.0 28.3 30.1 0.0 26.8
Function Calling Complete Enabled 0.0 0.0 42.9 16.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 58.3 18.8 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 49.6 7.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 48.1 4.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Orch-Isolated React Complete Enabled 2.3 13.8 35.0 77.8 11.0 63.2 68.1 96.8 24.4 22.1 0.0 36.2
Complete Disabled 4.4 19.8 31.0 86.8 16.5 62.5 66.2 97.5 23.3 28.7 0.2 31.8
Function Calling Complete Enabled 0.0 0.0 0.4 19.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.2 10.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 0.0 8.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 0.0 4.2 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
Single Agent React Complete Enabled 0.0 0.0 0.0 11.8 0.0 0.5 0.0 0.0 0.2 0.0 0.0 0.0
Complete Disabled 1.9 0.2 4.2 8.5 0.2 0.0 0.0 0.0 3.3 0.0 0.0 0.0
Function Calling Complete Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.0 0.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Summarized Disabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Table 4: Pass@1 Hallucination Rate ↓ Bolded scores indicate highest score (most hallucinations) per model on use case
3.2 Evaluation Metrics
Our primary evaluation metric is the Acceptable Score, which measures the percentage of records that satisfy all three critical success conditions simultaneously:
$\text{Acceptable}=\frac{|\{r\in R:C(r)\land A(r)\land O(r)\}|}{|R|}\times 100\%$
where $R$ is the set of all test records, and for each record $r$:
  • $C(r)$ indicates correct tool choice
  • $A(r)$ indicates correct tool arguments ($A(r)=1.0$)
  • $O(r)$ indicates correct final decision
We define our main acceptability metric in this way because it comes closest to representing success in a real world agentic system. Simply arriving at the expected decision may be insufficient in enterprise deployments. If the model arrives at the right choice, that a leave request should be submitted example, but then does not call the tool to submit the request or notify the employee, the task is not complete. Furthermore, if the model selects all the correct tools and makes the correct final decision but uses the wrong input values for a tool, that could lead to adverse outcomes such as a request being made for the wrong employee or a database being updated with a hallucinated value. For this reason, we consider an agentic flow to be successful if and only if all the expected tools have been called with expected values and the correct final decision is made based on the results of the tool calls.
Tool Choice Evaluation We evaluate tool choice under two conditions:
Lenient Acceptable ($C_{L}(r)$): The model executes all required ground truth tools but may perform additional read-only tools. Extraneous write operation tools are penalized.
Strict Acceptable ($C_{S}(r)$): The model must execute exactly the ground truth tools in the correct order with no extraneous tools and zero hallucinations.
Tool Arguments Scoring Tool arguments receive percent scoring where $A(r)=1.0$ only when all arguments are correct (compared to ground truth).
Reliability Metrics To assess system reliability, we evaluate each configuration over $k=8$ attempts and compute pass@1—the percent of correct samples over k trials) (DeepSeek-AI et al. 2025) and passˆK—the probability of all k trials succeeding (Yao et al. 2025). The latter metric is particularly important in real world enterprise domains where reliability and consistency is critical.
We use Acceptable pass@1 as our primary metric. We use lenient correct tool choice for the acceptable score as opposed to strict as strict scoring remains extremely challenging for most models. We use pass@1 as it brings stability to the lenient acceptable metric over k attempts, representing the success rate for a given model and configuration.
Additional Metrics We track supplementary behavioral metrics including:
  • Hallucination rate: Frequency of selecting non-existent agents or tools
  • Tool Repetition rate: Frequency of consecutive identical tool calls with identical arguments
4 Results
4.1 Overall Results
The benchmark results reveal that complex enterprise workflows remains a challenging task even for state-of-the-art language models in agentic systems. The simpler requesting time-off task (TO) showed had a peak score of 70.8% (GPT-4.1), while the complex customer routing task (CR) proved significantly more difficult, with scores peaking at only 35.3% (Sonnet 4). Only three models—GPT-4.1, Sonnet 4, and o3-mini—demonstrated meaningful capability on the complex customer request routing task, with most configurations achieving only single-digit scores. Performance varied dramatically across architectural configurations, indicating that no single architecture is optimal for all models. Furthermore, models achieved peak scores on different configurations between use cases suggesting that the optimal agentic configuration is in part dependent on the complexity of the task.
With regards to architecture elements, thinking tools consistently improved performance for non-reasoning models on the simpler time off task. GPT-4.1 improved from 48.5% to 70.8% with thinking tools enabled in single-agent function calling with summarized memory. However, thinking tools provided minimal benefit for o3-mini (55.8% to 56.7%) in its best performing category. Thinking tools showed minimal impact on the complex customer routing use case across all models. Compared to other elements, variation in memory management styles and orchestration strategies (within multi-agent architectures) both had minimal impact on scores.
Function calling generally outperformed ReAct across most models. The notable exception was LLaMA 3.3 70B, which achieved its peak score of 12.2% using single-agent ReAct on the requesting time off task. Multi-agent ReAct consistently underperformed suggesting that ReAct may be better suited towards single-agent systems. Hallucinations for all models (except for GPT-4o) were found exclusively under ReAct settings suggesting that while models may be robust to hallucinations on the function calling schemas they have been trained on, they still have the propensity to hallucinate in ReAct settings. This is particularly relevant on the more complex use case in multi-agent systems. Sonnet 4, for example, showed hallucination rates of 36-36% on ReAct in multi-agent systems but 0% hallucination in every other configuration. See Table 4 for hallucination rates.
While some models achieved their highest overall scores on single-agent architectures, the correct final decision scores for the complex customer routing use case were generally higher for multi-agent architectures. Although Sonnet 4’s highest overall scores were all achieved with single-agent function calling (all scores above 33%), its highest correct final decision scores were for function calling with multi-agent systems (84-87%) while for single agent it was in the 72-76% range. GPT-4.1 showed similar results as it achieved its peak overall score in both use cases with single agent function calling; however, on the customer routing task it achieved correct final decision rates of 97-99% with multi-agent function calling and only 79-86% with single agent function calling. GPT-4.1-mini also achieved significantly higher correct final decision rates with multi-agent function calling than with single-agent function calling on both use cases. This suggests that while some models are more likely to choose the correct tool or tool arguments on single agent systems, multi-agent systems may be more effective for selecting the correct final decision after all tool calls have been made. See Table 5 for more details.
Two key patterns emerged from the consistency analysis shown in Table 3. Sonnet 4 and GPT-4.1 showed the most robust performance with relatively low coefficients of variation (32.1% and 27.0% respectively for the simpler task), making them the most reliable choices across different agentic configurations. Conversely, o3-mini exhibited extreme sensitivity to architectural choices (CV = 143.7%).
Reliability Analysis. Passˆk scores across all models and agentic configurations peak at 0.0634 indicating only a 6.34% chance of executing the workflow correctly in all 8 trials. In enterprise settings, where reliability and consistency are essential, pass@1 and passˆk results reveal a fundamental gap between the promise of agentic LLMs and their real-world performance. See appendix A.4 for full scores.
4.2 Model Specific Results
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 71.0 75.0 14.6 30.2 10.6 4.0 15.6 5.8 1.7 0.0 87.9 68.0
Complete Disabled 70.8 73.2 13.3 26.2 10.6 3.8 13.8 3.8 1.0 0.0 86.5 65.2
Function Calling Complete Enabled 73.3 97.5 69.0 64.0 83.3 89.5 83.3 92.2 0.0 0.0 75.2 85.2
Complete Disabled 66.0 98.5 64.8 64.8 77.7 89.0 83.5 91.8 0.0 0.0 64.2 87.5
Summarized Enabled 74.8 99.2 69.8 64.8 79.8 86.8 83.8 89.0 0.0 0.0 76.0 86.8
Summarized Disabled 69.4 98.5 68.1 64.0 74.6 84.5 81.9 91.0 0.0 0.0 67.3 86.5
Orch-Isolated React Complete Enabled 71.9 83.5 11.5 33.0 7.3 4.5 15.8 4.2 0.4 0.0 88.3 63.7
Complete Disabled 72.7 82.0 13.8 31.2 8.5 5.0 14.4 4.0 0.8 0.0 88.5 70.5
Function Calling Complete Enabled 71.7 99.2 66.7 65.0 85.2 92.2 84.2 94.2 0.0 0.0 76.5 84.5
Complete Disabled 72.1 97.8 68.1 64.8 79.2 92.5 82.7 91.5 0.0 0.0 64.2 85.8
Summarized Enabled 79.2 97.8 67.3 64.5 83.8 90.0 82.9 92.5 0.0 0.0 74.0 86.8
Summarized Disabled 75.4 97.8 64.4 66.5 75.0 92.8 82.9 91.0 0.0 0.0 66.9 84.0
Single Agent React Complete Enabled 73.5 57.5 41.7 32.5 59.8 47.5 52.7 58.5 14.6 0.0 74.4 71.5
Complete Disabled 76.7 58.2 26.5 29.0 50.4 49.0 38.5 47.0 21.2 0.0 72.3 70.5
Function Calling Complete Enabled 72.3 86.5 64.2 66.2 59.4 23.5 83.8 18.0 0.0 0.0 76.7 75.5
Complete Disabled 69.4 79.2 62.1 60.5 42.9 21.5 84.8 18.0 17.9 0.0 44.8 72.5
Summarized Enabled 74.4 85.0 61.9 65.0 59.0 23.8 82.3 14.2 8.5 0.0 76.5 76.0
Summarized Disabled 70.4 80.8 58.8 61.8 46.0 22.8 83.3 14.2 0.0 0.0 44.6 74.0
Table 5: Pass@1 Correct Final Decision Rate. Bolded scores indicate highest score for a model on each use case.
Sonnet 4 achieved the highest scores on the complex customer routing task with a peak score of 35.3% using single-agent function calling. Sonnet 4 performed strongly on the requesting time off task (68.5% using orchestrator-isolated function calling). Notably, thinking tools had a substantial impact on the simpler task but minimal effect on the complex customer routing task. Time off scores improved from 64.4% to 68.5% with thinking tools, while customer routing performance remained relatively flat across configurations (33.3%-35.3% range).
GPT-4.1 achieved the highest overall score on the simpler use case - 70.8% using single-agent function calling with summarized memory and thinking tools. The model showed sensitivity to thinking tools in single-agent configurations, improving from 48.5% to 70.8% when enabled. However, thinking tools provided minimal benefit in multi-agent setups (54.4% vs 58.8%). Memory type preferences varied by orchestration: single-agent performance was nearly identical with complete (68.%1) versus summarized (70.8%) memory, while orchestrated configurations slightly favored complete memory.
o3-mini demonstrated extreme architectural sensitivity, achieving its peak score of 56.7% on the requesting time off use case with single-agent function calling but dropping to as low as 1.3% with orchestrated ReAct. Comparing enabled/disabled configurations showed negligble differences (e.g., 55.8% vs 56.7% in single-agent function calling). The model achieved its highest customer routing score of 22.3% using orchestrator-isolated function calling, substantially outperforming its single-agent CR performance (1.5%-2.0% range). This suggests that o3-mini is not inherently suited to single-agent vs multi-agent architecture but rather that its performance depends heavily on the complexity of the task.
GPT-4o and GPT-4.1-mini exhibited similar patterns, both peaking with orchestrator-isolated function calling configurations (53.5% and 67.1% TO respectively) using complete memory and thinking tools. Both models benefited substantially from thinking tool on simple tasks. GPT-4o improved from 26.9% to 39.0% in single-agent function calling with thinking tools, but they had minimal impact in orchestrated scenarios. GPT-4o performed extremely poorly with ReAct. GPT-4.1-mini, however, achieved higher scores on ReAct (41.5% and 51.2%) than on function calling (19.8-35.4%) within single-agent set ups on the time off use case. This finding illustrates the importance of testing agentic architectural elements in conjunction with each other since GPT-4.1-mini performed drastically differently on ReAct in multi-agent vs single-agent systems.
LLaMA 3.3 70B showed poor overall performance scoring 0 on most configurations. However, it achieved its highest score of 12.2% on the time off use case using single-agent ReAct without thinking tools, making it the only model to perform best with ReAct. LLaMA’s poor performance is likely due to its tool repetition rate which averaged 76.0% on the time off use case. Results on the customer routing case were almost entirely 0 due to persistent difficulties with instruction following and the function-calling format.
5 Discussion
We now use the above findings to make concrete recommendations for designing agentic AI systems in enterprise.
Avoiding Multi-Agent ReAct. One of the few universal trends that this benchmark shows is that models do not perform their best on Multi-Agent ReAct. In our set up, we used ReAct both for the orchestrator to select agents and for the agents to select tools. It is possible that using ReAct for just one instead of both of these components might improve outcomes but further work would be needed to test this.
Multi-Agent Systems for Final Decisions. While some models performed better overall on single-agent systems, we find that multi-agent systems are generally more effective at arriving at the correct final decision. For enterprise use cases where arriving at a final decision is more important than exact tool selection, we recommend the use of multi-agent systems.
Use Case Considerations. We observe that models tend to achieve peak performance under different architectures depending on the specific use case, rather than consistently performing best under a single architecture. As such, we caution developers against assuming a single architecture is particularly suited for any model and instead recommend testing against several use cases.
Using Thinking Tools. Our results indicate that for non-reasoning models, thinking tools can be a useful way for enabling more careful analysis on the information collected over the course of an agentic workflow. This is particularly true in contexts where the model must make a calculated decision based on the information it has collected (as is the case with the requesting time off use case). Thinking tools, however, can also add latency so we recommend developers consider carefully if their use cases require the extra reasoning abilities that these tools provide.
Model Selection. While large state-of-the-art models (in our benchmark GPT-4.1 and Claude Sonnet 4) perform the best across architectures, our benchmark shows that on simpler tasks, it is possible to achieve strong performances with smaller models if the appropriate architecture is chosen. For example, while GPT-4.1-mini’s scores on the time off use case dip as low as 4.6%, it’s peak acceptable score of 67.1% is comparable to Sonnet 4’s peak score of 68.5%.
6 Limitations
Our benchmark has several important limitations. We evaluate only two enterprise use cases with 60 samples each, which may not generalize across all industries and organizational contexts. We design one use case to represent a simple workflow and another as a more complex task, but a wider range of use cases would improve understanding of how task complexity affects performance on various architectures. Another key limitation is our model selection which is limited to six models, with only one open-source model (LLaMA 3.3 70B) and one reasoning-specialized model (o3-mini). Inclusion of more reasoning models in particular would provide more conclusive results on whether or not the results seen for o3-mini can be generalized to reasoning models as a whole. Finally, our use of deterministic tool responses enables controlled evaluation but may not reflect the variability of production enterprise systems.
7 Future Work
Several directions remain for expanding this benchmark to better capture enterprise agentic system requirements.
Multimodal Enterprise Integration. Our current evaluation is limited to text-based tool inputs and outputs, while real enterprise environments frequently involve multimodal data processing including document analysis, image recognition, and file processing workflows. Future work should expand the benchmark to include tools that accept and return images, PDFs, and other file formats commonly encountered in enterprise systems.
Conversational Capabilities Our benchmark intentionally focuses on autonomous workflows without user interaction as this is an important and understudied application of agentic, particularly in enterprise contexts. However, many enterprise applications require dynamic user communication and multi-turn conversations. Future evaluations should incorporate both static settings as well as conversational settings where agents must clarify requirements and adapt to changing user needs throughout task execution.
Multi-Agent ReAct. Our benchmark shows poor performances on multi-agent ReAct across models but does not fully illuminate the reason for this poor performance. Future work focusing specifically on Multi-Agent ReAct and testing ReAct for orchestration and tool calling independently in multi-agent systems would provide more insights on whether ReAct can be effectively utilized in multi-agent systems.
8 Conclusion
We introduce AgentArch, a benchmark for evaluating LLMs on a comprehensive agentic architectures on enterprise use cases. We test 18 configurations (single agent vs multi-agent, ReAct vs function calling, integration of thinking tools, and memory management styles) and find that even state-of-the-art LLMs struggle to maintain reliable performance on complex enterprise workflows in any set up. We also observe model-specific architectural preferences across LLMs suggesting that there is no universally optimal agentic architecture. Furthermore, even the best performing models do not demonstrate reliability over multiple trials showing a fundamental gap between the potential of agentic systems and their effectiveness in practice. We hope these findings provide actionable insights that enterprise developers can use to improve their model selection and architecture design of agentic systems.
Appendix A Appendix
A.1 Use Cases and Set Up
When evaluating single agent modes, we simply concatenate all the agent instructions for each defined agent and put them within tags to indicate which step the instructions refer to. In this way, agents in the single agent set up have access to all the same information as in multi-agent set ups.
Within each use case, we represent a wide array of potential edge cases. For example, in the Requesting Time Off use case we have user requests which should be rejected and some that should be accepted. Within the rejected category, we include requests that should be rejected for a wide variety of reasons such as invalid leave types, conflicting leaves, and most significantly, lack of leave balance. For lack of balance, we make sure to include complex cases where models must correctly count the number of days across multiple months or take into account leap years and the differing number of days in different months.
For each sample in a use case, the ground truth values for expected tool inputs, expected outcome, and expected tool order were human annotated and verified.
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 8.1 19.5 14.4 1.2 13.3 2.5 5.8 0.0 4.2 0.0 14.0 0.0
Complete Disabled 5.2 15.8 11.5 1.2 10.6 3.0 6.2 0.0 2.3 0.0 21.2 0.5
Function Calling Complete Enabled 52.3 17.5 27.9 18.0 14.2 16.2 2.3 18.2 0.0 0.0 58.8 16.0
Complete Disabled 49.6 21.5 17.1 19.8 12.1 16.8 2.1 19.8 0.0 0.0 49.8 17.2
Summarized Enabled 49.0 17.5 24.2 17.0 15.4 19.0 2.9 16.8 0.0 0.0 61.3 16.8
Summarized Disabled 49.6 17.5 24.0 16.8 6.2 21.5 2.7 17.0 0.0 0.0 57.1 16.5
Orch-Isolated React Complete Enabled 10.2 16.0 8.3 0.2 12.1 3.2 7.3 0.0 6.0 0.0 12.1 0.0
Complete Disabled 9.8 15.2 9.8 0.0 10.4 1.2 5.0 0.0 3.3 0.0 18.8 0.0
Function Calling Complete Enabled 50.0 17.5 47.1 13.8 6.0 0.8 2.5 20.2 0.0 0.0 61.0 15.0
Complete Disabled 50.2 17.0 50.8 12.2 6.5 1.5 1.5 18.5 0.0 0.0 55.0 16.2
Summarized Enabled 44.4 16.5 39.0 12.8 16.9 2.0 0.8 18.5 0.0 0.0 62.3 16.8
Summarized Disabled 44.4 16.0 43.8 14.2 10.8 1.5 2.3 19.5 0.0 0.0 57.9 15.0
Single Agent React Complete Enabled 30.0 5.2 20.8 1.8 0.2 7.2 8.8 1.5 0.8 0.0 58.8 17.2
Complete Disabled 39.4 9.8 13.1 1.0 0.0 5.2 4.0 2.2 0.8 0.0 47.9 31.0
Function Calling Complete Enabled 68.5 12.2 31.9 4.2 0.0 0.2 38.5 1.2 0.0 0.0 3.3 37.2
Complete Disabled 52.3 18.5 26.7 4.2 0.2 0.2 37.7 2.5 0.0 0.0 0.0 36.0
Summarized Enabled 71.7 13.2 30.2 3.2 0.0 0.8 36.5 1.8 0.0 0.0 3.1 35.8
Summarized Disabled 49.0 17.2 23.5 3.2 0.0 0.2 39.2 1.2 0.0 0.0 0.0 36.2
Table 6: Pass@1 Correct Tools Strict Rate. Bolded scores indicate highest score per model on use case
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 37.9 21.8 26.5 7.0 32.3 19.5 8.8 0.2 16.2 4.5 28.1 30.8
Complete Disabled 32.1 19.0 23.1 10.2 24.4 26.0 9.2 0.0 12.7 6.6 29.0 43.2
Function Calling Complete Enabled 58.8 25.8 51.2 24.2 69.2 43.8 4.6 20.5 13.3 0.0 68.3 44.0
Complete Disabled 54.8 28.5 46.5 26.8 56.9 51.7 5.4 21.8 14.2 0.0 57.7 46.0
Summarized Enabled 55.6 23.2 55.4 25.5 68.3 42.2 5.2 18.5 0.0 0.0 69.8 42.0
Summarized Disabled 55.2 25.2 51.2 24.2 54.8 46.8 5.2 19.2 0.0 0.0 57.9 43.2
Orch-Isolated React Complete Enabled 29.2 17.8 25.0 1.0 27.1 8.8 9.4 0.5 18.1 2.0 28.5 3.2
Complete Disabled 26.2 17.0 24.2 1.8 26.9 7.5 8.3 0.2 13.5 0.8 28.1 2.2
Function Calling Complete Enabled 58.8 17.8 53.5 14.0 70.4 0.8 6.0 22.2 3.1 0.0 70.2 42.8
Complete Disabled 55.0 17.0 52.3 13.0 58.5 1.8 4.0 21.0 0.6 0.0 57.5 16.2
Summarized Enabled 55.8 16.5 50.6 14.0 69.0 2.0 2.9 20.5 0.0 0.0 66.9 16.8
Summarized Disabled 53.8 16.0 49.0 15.2 60.0 1.5 5.0 21.2 0.0 0.0 57.9 15.0
Single Agent React Complete Enabled 36.5 8.0 25.4 2.5 57.9 19.2 9.2 1.8 11.2 4.8 63.7 30.0
Complete Disabled 46.0 16.0 19.8 3.5 59.4 13.0 4.2 2.5 12.7 7.9 54.2 35.8
Function Calling Complete Enabled 69.0 29.0 40.6 5.5 43.5 12.2 56.9 1.5 0.0 0.0 66.7 43.5
Complete Disabled 55.2 36.5 29.4 7.2 37.5 12.8 56.0 3.0 5.0 0.0 31.7 50.7
Summarized Enabled 71.9 29.5 40.2 6.5 41.9 13.0 55.0 2.8 0.8 0.0 64.4 42.0
Summarized Disabled 53.3 34.5 30.8 6.0 41.5 16.0 56.0 1.2 0.0 0.0 31.7 49.0
Table 7: Pass@1 Correct Tools Lenient Rate. Bolded scores indicate highest score per model on use case
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Function Calling Complete Enabled 1.4 0.0 0.4 0.0 3.5 0.0 0.0 0.0 0.0 0.0 4.4 0.0
Complete Disabled 0.8 0.0 0.2 0.0 1.0 0.0 0.0 0.0 0.0 0.0 1.2 0.0
Summarized Enabled 0.5 0.0 0.3 0.0 1.5 0.0 0.0 0.0 0.0 0.0 1.6 0.0
Summarized Disabled 0.6 0.0 0.2 0.0 0.4 0.0 0.0 0.0 0.0 0.0 0.1 0.0
Orch-Isolated React Complete Enabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Complete Disabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
Function Calling Complete Enabled 1.4 0.0 0.7 0.0 4.1 0.0 0.0 0.0 0.0 0.0 4.9 0.0
Complete Disabled 0.8 0.0 0.5 0.0 1.1 0.0 0.0 0.0 0.0 0.0 1.2 0.0
Summarized Enabled 0.8 0.0 0.2 0.0 1.6 0.0 0.0 0.0 0.0 0.0 1.3 0.0
Summarized Disabled 0.5 0.0 0.1 0.0 0.8 0.0 0.0 0.0 0.0 0.0 0.2 0.0
Single Agent React Complete Enabled 0.0 0.0 0.0 0.0 0.5 0.0 0.0 0.0 0.0 0.0 2.7 0.0
Complete Disabled 0.0 0.0 0.0 0.0 0.1 0.0 0.0 0.0 0.0 0.0 0.6 0.0
Function Calling Complete Enabled 4.6 0.0 0.1 0.0 0.0 0.0 1.1 0.0 0.0 0.0 3.9 0.0
Complete Disabled 0.4 0.0 0.0 0.0 0.0 0.0 0.9 0.0 0.0 0.0 0.0 0.0
Summarized Enabled 6.3 0.0 0.0 0.0 0.0 0.0 0.8 0.0 0.0 0.0 2.9 0.0
Summarized Disabled 0.3 0.0 0.0 0.0 0.0 0.0 0.9 0.0 0.0 0.0 0.0 0.0
Table 8: Passˆ K Acceptable Rate. Bolded scores indicate highest score per model on use case
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 7.5 28.2 57.7 36.2 55.2 93.5 64.2 1.5 90.0 84.2 0.6 56.8
Complete Disabled 9.4 35.2 58.1 41.8 53.8 94.0 64.2 3.2 94.6 96.8 0.0 40.2
Function Calling Complete Enabled 9.4 0.0 9.4 5.0 9.6 17.5 0.0 0.0 100.0 100.0 0.0 0.0
Complete Disabled 11.0 0.0 10.6 3.8 15.6 16.8 0.0 0.0 100.0 100.0 0.0 0.0
Summarized Enabled 12.5 0.0 3.8 2.8 24.0 19.5 0.0 0.0 100.0 100.0 0.0 0.0
Summarized Disabled 13.1 0.0 4.2 1.8 31.0 18.8 0.0 0.0 100.0 100.0 0.0 0.2
Orch-Isolated React Complete Enabled 10.6 17.2 54.0 20.0 54.6 90.0 62.1 1.0 91.5 92.4 1.5 26.0
Complete Disabled 10.2 22.8 46.2 28.7 61.3 81.8 62.7 2.0 95.6 91.2 0.4 24.2
Function Calling Complete Enabled 6.9 0.2 1.7 7.8 9.8 18.2 0.0 0.0 100.0 100.0 0.0 0.0
Complete Disabled 12.3 0.0 1.5 2.8 21.2 13.2 0.0 0.0 100.0 100.0 0.0 0.0
Summarized Enabled 9.0 0.2 4.4 5.5 24.8 17.0 0.0 0.0 100.0 100.0 0.0 0.0
Summarized Disabled 16.0 0.0 2.3 6.5 32.5 13.0 0.0 0.0 100.0 100.0 0.0 0.0
Single Agent React Complete Enabled 1.0 22.8 4.4 3.2 20.4 46.5 0.0 0.0 0.0 24.2 0.0 8.8
Complete Disabled 1.9 21.2 2.9 5.5 9.4 29.8 0.0 0.0 0.4 12.7 0.0 8.0
Function Calling Complete Enabled 0.0 5.8 0.2 6.0 0.2 4.5 0.0 0.0 0.0 88.9 0.0 0.0
Complete Disabled 0.4 2.8 0.0 5.0 1.5 6.5 0.0 100.0 98.8 100.0 0.0 0.0
Summarized Enabled 0.0 4.2 0.6 5.2 0.6 3.5 0.0 100.0 99.2 100.0 0.0 0.0
Summarized Disabled 0.4 3.2 0.0 4.8 1.7 5.0 0.0 0.0 0.0 98.8 0.0 0.0
Table 9: Pass@1 Tool Repetition Rate ↓. Bolded scores indicate highest score (most repetition) per model on use case
Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 36.0 65.5 32.3 62.0 32.3 58.8 74.6 63.2 31.0 68.2 3.1 15.5
Complete Disabled 43.5 65.8 34.6 60.8 36.9 50.5 70.4 62.3 41.9 85.4 3.3 5.8
Function Calling Complete Enabled 13.8 58.5 5.0 56.2 6.2 27.3 94.6 74.2 85.8 100.0 0.6 40.0
Complete Disabled 9.0 56.5 2.7 54.2 5.0 16.2 94.0 73.8 83.5 100.0 3.8 40.0
Summarized Enabled 7.1 61.3 4.2 54.8 5.4 27.8 94.2 74.2 100.0 100.0 2.1 40.0
Summarized Disabled 6.5 60.0 3.3 55.2 6.9 18.2 94.0 73.5 100.0 100.0 1.9 40.0
Orch-Isolated React Complete Enabled 55.2 70.0 40.0 72.8 43.8 70.2 76.2 63.7 28.5 69.2 2.3 66.8
Complete Disabled 57.5 69.8 42.5 72.8 39.0 70.5 73.3 62.3 42.1 77.8 2.7 68.2
Function Calling Complete Enabled 13.5 69.8 10.8 69.8 5.0 70.0 92.5 73.0 96.9 96.4 0.6 40.2
Complete Disabled 16.7 70.0 1.9 69.2 4.4 70.0 95.0 72.8 99.4 98.2 2.5 70.0
Summarized Enabled 19.4 70.0 11.2 68.8 4.0 70.0 95.6 73.0 100.0 98.8 3.1 70.0
Summarized Disabled 17.3 70.0 5.0 69.0 2.9 70.0 94.2 74.0 100.0 100.0 3.5 70.0
Single Agent React Complete Enabled 34.0 62.0 40.8 74.0 23.5 57.0 44.6 83.5 47.1 79.2 10.6 33.8
Complete Disabled 32.9 55.2 46.7 75.8 28.3 62.0 34.8 77.8 51.9 82.4 16.7 28.5
Function Calling Complete Enabled 3.5 41.0 18.8 65.0 16.9 57.8 24.8 94.8 100.0 100.0 10.2 22.5
Complete Disabled 16.5 33.5 24.6 62.7 18.8 57.8 26.2 92.5 91.9 100.0 11.5 19.5
Summarized Enabled 3.1 40.5 18.3 65.0 16.5 56.5 25.4 93.5 98.3 100.0 12.5 24.2
Summarized Disabled 17.3 35.5 18.3 64.8 15.4 53.5 24.8 94.2 100.0 100.0 11.7 20.2
Table 10: Pass@1 Missing Required Tool Rate ↓. Bolded scores indicate highest score (most missing tools) per model on use case
A.2 Thinking Tools Examples
Below are examples of how thinking tools can be used by the model to complete tasks in agentic frameworks. Thinking tools effectively give the model space to generate extra tokens in the format of calling a tool. There is no actual tool that gets executed when these tools are selected. Whatever the model passes as the tool argument is returned as the result and added to memory so the model can use the information later on in the task.
  1. 1.
    Synthesizing information
    tool: synthesize_collected_information
    args:
    synthesized_information (str):
    "I need to determine whether or not the employee Sarah Smith is eligible for leave. I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days. Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
    result (same as input):
    "I need to determine whether or not the employee Sarah Smith is eligible for leave. I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days. Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
  2. 2.
    Math
    tool: math
    args:
    input (str):
    "Sarah Smith requested leave from ’all of February 2024 to the second of March’. I need to figure out how many days that is. Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
    result (same as input)::
    "Sarah Smith requested leave from ’all of February 2024 to the second of March’. I need to figure out how many days that is. Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
A.3 Prompts
Prompts were designed to be simple and representative examples of both traditional function calling-style prompts and ReAct prompts. For function calling prompts, the tools and directions are placed in the system prompt while the user prompt includes the main objective and conversation history thus far. We also include a note to always call a tool (since communication with the user is not supported), always use real input values (since some models are prone to using placeholders instead), and to select the finish tool when finished since this is critical for control flow as well as getting a self reported summary of each agents actions which is passed to other agents in the summarized memory management style. We attempt to convey the same information in the ReAct style prompt however as is typical with ReAct prompting, all information is in the system prompt and a JSON response is required. The ReAct prompts include a sample JSON output format that asks the model for ”thought” (reasoning) and an ”action” and ”observation.”.
A.4 Additional Metrics
We include supplemental metrics here all shown with pass@1: Correct Tools Strict Rate, Correct Tools Lenient Rate, Tool Repetition Rate (higher scores indicate more repetition), Missing Required Tool Rate (higher scores indicate more missing tools). We also show PassˆK for the Acceptable Rate.
system:
You are an agent. Your goal is to accomplish your main objective using the directions provided. You must follow the directions you are given. If the directions instruct you to complete steps in a specific order, you must follow that order.
Directions: {agent_instructions}
user:
Main Objective: {main_objective}
Conversation History: {memory}
Always call a tool. You must always pass real values as tool inputs. You may never use placeholders. Before calling a tool, remember to consider the conversation history to understand what has already been accomplished and what needs to be done next. When you have finished your task, you MUST select the finish tool. Otherwise your response will be ignored.
Figure 3: Function Calling Agent Prompt
user: You are an agent. Your goal is to accomplish your main objective using the directions provided. You must follow the directions you are given. If the directions instruct you to complete steps in a specific order, you must follow that order.
Directions:
{agent_instructions}
Main Objective:
{main_objective}
Conversation History:
{memory}
To accomplish your main objective, you must select an action from the following options:
{tools}
Always select an action using the format below. You must always pass real values as arguments. You may never use placeholders. Before choosing an action, remember to consider the conversation history to understand what has already been accomplished and what needs to be done next. When you have finished your task, you MUST select the finish action. Otherwise your response will be ignored.
Output your response in the following JSON format:
"thought": "<reason step by step about the current situation and what to do next>",
"action": {
"name": "<tool_name_1>",
"arguments": {
"<argument_name>": <argument_value>
"observation": "<expected result of the tool execution>"
"thought": "<reason step by step about the current situation and what to do next>",
"action": {
"name": "<tool_name_m>",
"arguments": {
"<argument_name>": <argument_value>
"observation": "<expected result of the tool execution>"
Figure 4: ReAct Agent Prompt
user: You are the orchestrator of a group of agents. Your job is to select the next agent(s) best suited for accomplishing the request given the conversation history thus far. You must follow the directions you are given. If the directions instruct you to complete steps in a specific order, you must follow that order.
The conversation history below shows you what agents have already done to accomplish the task.
Directions:
{usecase_directions}
Request:
{request}
Conversation History:
{memory}
To accomplish your request, you must select from the following agents:
{agents}
Always select an agent using the format below. You must always pass real values as arguments. You may never use placeholders. Before selecting the next agent(s), consider your request, directions, and the conversation history thus far to understand what to do next.
Output your response in the following JSON format:
"thought": "<reason step by step about the current situation and what to do next>",
"action": {
"name": "<agent_name_1>",
"arguments": {
"main_objective": <main objective you want the agent to accomplish>
"observation": "<expected result of the agent execution>"
"thought": "<reason step by step about the current situation and what to do next>",
"action": {
"name": "<agent_name_n>",
"arguments": {
"main_objective": <main objective you want the agent to accomplish>
"observation": "<expected result of the agent execution>"
If there are no relevant agents, respond with "Unfortunately no agents are able to help with this request." Otherwise you MUST respond with at least one agent in the format above. .
Figure 5: Orchestrator ReAct Prompt
system:
You are the orchestrator of a group of agents. Your job is to select the next agent(s) best suited for accomplishing the request. You must follow the directions you are given. If the directions instruct you to complete steps in a specific order, you must follow that order. When you select an agent, you must give it a main objective that describes what the agent should accomplish.
You also have access to the conversation history below which shows you what agents have already done to accomplish the task.
Directions: {usecase_directions}
user:
Request: {user_utterance}
Conversation History: {memory}
Always select an agent. You must always pass real values as arguments. You may never use placeholders. Before selecting the next agent(s), consider your request, directions, and the conversation history thus far to understand what to do next. If there are no relevant agents, respond with "Unfortunately no agents are able to help with this request." Otherwise you MUST respond with at least one agent .
Figure 6: Orchestrator ReAct Prompt

### v2

@@
-AgentArch: A Comprehensive Benchmark to Evaluate Agent Architectures in Enterprise
+AgentArch: A Benchmark for Evaluating Agent Architectures in Enterprise Workflows
 Abstract
-While individual components of agentic architectures have been studied in isolation, there remains limited empirical understanding of how different design dimensions interact within complex multi-agent systems.
-This study aims to address these gaps by providing a comprehensive enterprise-specific benchmark evaluating 18 distinct agentic configurations across state-of-the-art large language models.
-We examine four critical agentic system dimensions: orchestration strategy, agent prompt implementation (ReAct versus function calling), memory architecture, and thinking tool integration.
-Our benchmark reveals significant model-specific architectural preferences that challenge the prevalent one-size-fits-all paradigm in agentic AI systems.
-It also reveals significant weaknesses in overall agentic performance on enterprise tasks with the highest scoring models achieving a maximum of only 35.3% success on the more complex task and 70.8% on the simpler task.
-We hope these findings inform the design of future agentic systems by enabling more empirically backed decisions regarding architectural components and model selection.
-[^1] GitHub Repository: https://github.com/ServiceNow/AgentArch
+While individual components of agentic architectures have been studied in isolation, we lack empirical evidence on how design dimensions interact in realistic multi-agent systems.
+This study aims to address these gaps by providing a comprehensive enterprise-specific benchmark.
+We introduce AgentArch—a benchmark that evaluates 18 agentic configurations across state-of-the-art large language models.
+The study spans four dimensions: orchestration strategy, agent style (ReAct versus function calling), memory architecture, and thinking tool integration.
+Results reveal strong model-specific architectural preferences that challenge one-size-fits-all assumptions in agentic AI.
+They also expose notable performance gaps on enterprise tasks: the best models reach only 35.3% success on the complex task and 70.8% on the simpler task.
+We aim to inform architecting choices for enterprise agents by grounding model selection and component design in systematic evidence.
+AgentArch: A Benchmark for Evaluating Agent Architectures in Enterprise Workflows
+Tara Bogavelli, Hari Subramani, Roshnee Sharma ServiceNow Correspondence: tara.bogavelli@servicenow.com https://github.com/ServiceNow/AgentArch
 1 Introduction
-Large language models (LLMs) have evolved beyond simple text generation to function as autonomous agents capable of iterative decision-making and complex task completion (Yao et al.
-2025).
+Large language models (LLMs) have evolved beyond simple text generation to function as autonomous agents capable of iterative decision-making and complex task completion taubench:25.
 This transformation has opened new possibilities for enterprise automation, where agentic AI systems can interact with existing workflows, databases, and business processes to accomplish multi-step objectives with minimal human intervention.
@@
 The fragmented approach of prior benchmarks leaves enterprise developers without empirical guidance for selecting optimal combinations of architectural components or choosing appropriate models for their specific setups.
-This paper presents a comprehensive evaluation of 18 distinct agentic architectures across state-of-the-art LLMs, systematically examining the interactions between four key architectural dimensions: orchestration strategy (single-agent vs.
+This paper presents a comprehensive evaluation of 18 distinct agentic architectures across state-of-the-art LLMs, systematically examining the interactions between four key architectural dimensions: orchestration strategy (single agent vs.
 multi-agent), agent style (function calling vs.
 ReAct), memory management (complete vs.
-summary), and thinking tool integration (Yao et al.
-2023).
+summary), and thinking tool integration react:23.
 We test each architectural combination on two enterprise use cases of varying complexity, providing practical insights for agentic system designers.
@@
 Our results provide empirically-grounded guidance for practitioners building enterprise agentic systems, enabling more informed decisions about architectural components and model selection based on specific use case requirements.
-2 Related Work
+To summarize, our work makes the following contributions:
+  • We introduce AgentArch, the first benchmark systematically evaluating 18 agentic architectures across six large language models on realistic enterprise workflows.
+  • We jointly analyze the effects of orchestration, prompting style, memory design, and thinking tool integration, providing quantitative insight into how these factors interact in practice.
+  • Our experiments reveal model-specific architectural preferences, showing that optimal configurations vary by model and task complexity, rather than following a single best-performing design.
+  • We offer evidence-based recommendations for selecting models and architectural configurations suited to enterprise reliability and integration constraints.
+Overall, AgentArch contributes a systematic, empirically grounded step toward understanding how architectural design choices influence agentic system behavior, supporting more informed and reproducible development practices.
 Requesting Time Off (TO) Customer Request Routing (CR)
-Description Simple, structured workflow for PTO eligibility verification and request processing.
-Tests basic multi-step reasoning with clear success criteria.
-Complex workflow mimicking intelligent customer service systems.
-Handles simple requests automatically while escalating complex issues.
-Tests classification and escalation decisions.
+Description Simple workflow for PTO eligibility and request processing.
+Tests basic multi-step reasoning and counting.
+Complex workflow mimicking customer service systems.
+Automates simple requests, escalating complex issues.
+Tests classification and instruction following.
 Tools 8 custom tools 31 custom tools
-Agents 3 agents 9 agents
-Agent Types Employee Information Retriever Agent, Balance Eligibility Agent, Leave Approval Agent Request Validation Agent, Duplicate Case Detection Agent, Case Creation Agent, Informational Queries Agent, Transactional Queries Agent, Intent and Sentiment Extraction Agent, Entity Extraction Agent, Document Verification Agent, Email Generation Agent
+Agents Employee Information Retriever, Balance Eligibility, Leave Approval Request Validation, Duplicate Case Detection, Case Creation, Informational Queries, Transactional Queries, Intent and Sentiment Extraction, Entity Extraction, Document Verification, Email Generation
 Key Challenges Date calculations; leave balance verification; policy compliance; multi-step approval workflow Appropriate escalation decisions; context preservation; handling ambiguous requests; complex routing logic; instruction following
 Table 1: Detailed comparison of the two enterprise use cases evaluated in our benchmark.
-The evaluation of agentic AI systems has emerged as a critical research area, with existing benchmarks examining various components of agent architectures.
-However, these efforts have predominantly focused on isolated aspects of agentic systems rather than their integrated performance.
-Tool calling is perhaps the most fundamental aspect of all agentic systems.
-Existing tool calling benchmarks (Liu et al.
-2024b; Patil et al.
-2025; Wu et al.
-2024) have revealed significant LLM weaknesses in key aspects of tool calling including tool selection accuracy, tool input accuracy, valid output formatting, and issues with latency and unnecessary repetition of tool calls.
-Additional benchmarks focus on more complex scenarios finding that LLM performance on tool calling decreases when combined with complex instruction following (Yao et al.
-2025) and multi step reasoning in nested tool calling scenarios (Basu et al.
-2025).
-Another key aspect of agentic systems is memory and context management.
-LLMs are inherently limited by their context windows which provides a unique challenge for those building agentic systems to ensure that the LLMs keep the most relevant information in context.
-This is relevant for both conversational and tool-use settings, as conversational context accumulates over time, while tool calls can significantly increase context length due to their frequency or output size (Maharana et al.
-2024a).
-Existing research presents a variety of mechanisms for improving memory management in agentic systems such as treating agent memory as a self directed operating system with different storage tiers (Packer et al.
-2024) , creating dynamic knowledge networks for connecting information (Xu et al.
-2025), and implementing a dual-track memory system meant to resemble long-term and short-term memory in the human brain (Liu et al.
-2024a).
-As LLMs enable more complex agentic workflows, Multi-Agent Systems (MAS) have emerged to support use cases involving specialized agents, prompting the development of benchmarks to evaluate their effectiveness.
-(Zhu et al.
-2025) examine LLM performance across various MAS structures, including hierarchical and distributed planning, evaluating task completion and planning capabilities to assess LLM effectiveness in both planning and execution roles.
-(Agashe, Fan, and Wang 2025) and (Geng and Chang 2025) focus specifically on evaluating planning and coordination between agents in MAS finding that planning abilities are a key component of success in MAS.
-(Yue et al.
-2025) illustrate how multi-agent systems can benefit from routing frameworks that help ensure that the appropriate agents, communication strategies, and LLMs are selected for a given task.
-Similarly, (Hou, Tang, and Wang 2025) introduce a dynamic MAS framework that incorporates hierarchical planning along with a search process for selecting the optimal multi-agent configuration for the task at hand.
-Other work on dynamically generated MAS demonstrates how scaling single agents into MAS can lead to performance increases (Yuan et al.
-2025).
-Recent work compares heterogeneous LLMs versus a single model in MAS, highlighting how MAS can be used to leverage individual model strengths (Ye et al.
-2025).
-As our work is primarily focused on enterprise benchmarks, we also explore existing work on agents in enterprise domains.
-Enterprise-focused evaluation has consistently revealed a gap between existing academic benchmarks and real-world applicability, with even state-of-the-art models struggling on simple workplace tasks (Styles et al.
-2024).
-Benchmarks across different enterprise domains consistently show that realistic business task complexity significantly exceeds what current models can handle reliably, with performance typically degrading in multi-turn interactions and interconnected tasks (Huang et al.
-2025a, b; Drouin et al.
-2024; Styles et al.
-2024).
-While prior benchmarks have advanced understanding of individual agentic components, they remain fragmented—evaluating tool use (Yao et al.
-2025; Patil et al.
-2025) (Basu et al.
-2025), memory (Maharana et al.
-2024b), and coordination (Geng and Chang 2025; Agashe, Fan, and Wang 2025) in isolation.
-Existing enterprise benchmarks address task complexity but lack systematic analysis of how architectural choices interact across scenarios.
-As a result, developers lack empirical guidance for selecting and combining agentic components under real-world enterprise constraints.
-Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
-TO CR TO CR TO CR TO CR TO CR TO CR
-Orch-Open React Complete Enabled 36.9 21.2 6.2 0.0 7.1 0.0 1.2 0.0 0.4 0.0 28.1 0.0
-Complete Disabled 31.9 17.5 4.8 1.2 7.3 0.5 1.2 0.0 0.0 0.0 28.7 0.2
-Function Calling Complete Enabled 58.8 14.8 50.2 4.2 65.8 4.8 4.6 20.5 0.0 0.0 67.7 15.2
-Complete Disabled 54.4 16.2 46.5 5.0 56.0 4.8 5.4 21.8 0.0 0.0 57.7 17.2
-Summarized Enabled 51.9 15.2 49.0 0.8 59.0 2.5 5.0 18.5 0.0 0.0 59.6 16.8
-Summarized Disabled 52.7 15.8 45.0 0.8 49.6 1.0 4.8 19.0 0.0 0.0 44.2 16.5
-Orch-Isolated React Complete Enabled 28.7 17.2 3.8 0.0 4.6 0.0 2.3 0.0 0.2 0.0 28.3 0.0
-Complete Disabled 25.8 16.0 5.8 0.2 5.6 0.0 1.5 0.0 0.0 0.0 28.1 0.5
-Function Calling Complete Enabled 58.8 17.2 53.5 1.8 67.1 0.8 6.0 22.2 0.0 0.0 68.5 14.8
-Complete Disabled 54.6 16.2 52.1 1.8 57.1 1.5 4.0 21.0 0.0 0.0 57.5 16.2
-Summarized Enabled 54.6 16.5 46.0 1.8 59.6 2.0 2.9 20.5 0.0 0.0 58.1 16.8
-Summarized Disabled 51.5 15.8 44.2 3.0 54.4 1.5 4.8 21.2 0.0 0.0 44.8 14.5
-Single Agent React Complete Enabled 34.4 1.0 15.8 0.2 51.2 1.2 8.8 1.5 6.2 0.0 63.7 3.5
-Complete Disabled 36.7 3.8 9.6 0.0 41.5 1.2 2.9 2.2 12.3 0.0 52.7 9.0
-Function Calling Complete Enabled 68.1 22.2 39.0 3.0 35.4 0.2 56.7 1.5 0.0 0.0 66.7 33.8
-Complete Disabled 49.6 21.2 26.9 3.5 19.8 0.2 55.8 2.0 3.8 0.0 31.7 35.0
-Summarized Enabled 70.8 21.5 37.9 2.5 33.5 0.5 55.0 1.5 0.6 0.0 64.4 33.2
-Summarized Disabled 48.5 19.8 28.7 2.8 24.4 0.5 55.8 1.0 0.0 0.0 31.7 35.2
-Table 2: Pass@1 Acceptable Rate. Bolded scores indicate highest score per model on use case.
+2 Related Work
+Evaluating agentic AI systems has become a central research focus, yet many studies still assess components in isolation rather than integrated end-to-end performance.
+Tool calling.
+Tool use is foundational for agentic systems.
+Benchmarks such as agentbench:24; bfcl:25; sealtools:24 document persistent weaknesses in tool selection, argument fidelity, output formatting, latency, and redundant calls.
+More complex settings further degrade performance: combining tool use with intricate instruction following taubench:25 or multi-step, nested tool invocations basu2025nestfulbenchmarkevaluatingllms increases error rates.
+Memory and context management.
+Limited context windows and growing conversational and tool-response histories create pressure on memory design.
+Prior work explores tiered memory as an OS abstraction packer2024memgptllmsoperatingsystems, dynamic knowledge graphs for retrieval and linking xu2025amemagenticmemoryllm, and dual-track (short/long-term) schemes liu2024llmconversationalagentmemory.
+Empirical studies highlight how tool calls inflate context via frequent, verbose outputs maharana2024evaluatinglongtermconversationalmemory, motivating mechanisms that preserve relevance while controlling prompt length.
+Multi-Agent Systems (MAS).
+With increasing workflow complexity, MAS introduce specialization and coordination.
+Benchmarks examine planning structures (hierarchical vs.
+distributed) and execution quality marble:25, as well as inter-agent planning/coordination llmcoord:25; geng2025realmbenchrealworldplanningbenchmark, identifying planning reliability as a bottleneck.
+Routing frameworks improve agent/model selection and communication patterns masrouter:25; adaptive approaches search multi-agent configurations online hou2025halohierarchicalautonomouslogicoriented.
+Automatic scaling from single agents to MAS can yield gains yuan2025evoagentautomaticmultiagentgeneration, and heterogeneous-model MAS leverage complementary strengths ye2025xmasbuildingmultiagentsystems.
+Enterprise evaluations.
+Enterprise-focused studies consistently reveal a gap between academic benchmarks and real deployments.
+Even state-of-the-art models struggle on routine workplace tasks workbench:24, with performance degrading in multi-turn and interdependent workflows crmarena:25; huang2025crmarenaproholisticassessmentllm; workarena:24; workbench:24.
+These findings underscore the need for evaluations that reflect realistic business constraints: reliability, strict procedure orderings, and integration with existing systems.
+Prior work has advanced understanding of individual components—tool use taubench:25; bfcl:25; basu2025nestfulbenchmarkevaluatingllms, memory locomo:24, and coordination geng2025realmbenchrealworldplanningbenchmark; llmcoord:25—but remains fragmented.
+Enterprise-oriented benchmarks capture task complexity yet rarely analyze interactions among orchestration, prompting style, memory strategy, and reasoning aids.
+Our benchmark fills this gap by jointly varying these dimensions across two enterprise use cases, providing evidence on model-specific architectural preferences and the relative advantages of multi-agent versus single agent designs.
 Model TO CR
@@
 3.1 Experiment Set Up
-Use Cases We evaluate two enterprise workflows of varying complexity to examine how task complexity affects agentic performance on real-world enterprise automation scenarios.
-Both of these use cases involve specific directions to complete tasks in a specified order, reflecting the rigid nature of many enterprise workflows where steps must take place in a predefined order.
+Use Cases We evaluate two enterprise workflows of differing complexity to examine how task complexity affects agentic performance on real-world enterprise automation scenarios.
+Both of these use cases involve highly-specific directions to complete tasks in a given order, reflecting the rigid nature of many enterprise workflows where steps must take place in a predefined order.
 For each request in a use case, we define an expected final decision (e.g.
-leave approve or rejected).
+leave approved or rejected).
 We create mock data so that each tool request returns a deterministic result.
@@
 This approach better represents the information processing challenges agents face in production environments where data is rarely clean or perfectly structured.
-Figure 1: Orchestration, Isolated Agents. Agents ask orchestrator for help and orchestrator selects agent to assist.
-Figure 2: Orchestration, Open Agent Network. Agents ask other agents for help directly.
 Requesting Time Off (TO) represents a simple workflow involving PTO eligibility verification and request processing.
@@
 Each use case includes 60 user utterances designed to cover the full spectrum of real-world requests, including successful scenarios, edge cases, and failure conditions that would occur in actual enterprise deployments.
-Implementation details are provided in Appendix A.1.
+Implementation details are provided in Appendix A.
 Orchestration Strategies.
-We evaluate four orchestration strategies to examine single-agent versus multi-agent performance across key dimensions: decision-making with varying option sets, collaboration costs and benefits, and specialist versus generalist agent effectiveness.
+We evaluate three orchestration strategies to examine single agent versus multi-agent performance across key dimensions: decision-making with varying option sets, collaboration costs and benefits, and specialist versus generalist agent effectiveness.
 1) Orchestrator-led, Isolated Agent.
@@
 A single agent receives the user request and has access to all available tools, with no collaboration mechanisms.
-This tests agent performance in large, unfiltered option spaces.
+This tests agent performance in larger, unfiltered option spaces.
+Figure 1: Orchestration, Isolated Agents. Agents ask the orchestrator for help, and the orchestrator selects agent to assist.
 Agents.
@@
 With function calling, the model directly selects tools from the available toolkit, leveraging inherent function calling capabilities without explicit intermediate reasoning.
-Using ReAct, The model is instructed to output its thoughts before selecting actions thus using a structured reasoning-action framework (Yao et al.
-2023).
+Using ReAct, the model outputs its reasoning before selecting actions, following a structured reasoning-action framework react:23.
 We investigate whether the explicit reasoning in ReAct prompting yields substantially different outcomes compared to direct function calling.
@@
 Memory Management.
-We evaluate two information sharing approaches to examine the trade-off between context completeness and input context length.
+We evaluate two information-sharing approaches to examine the trade-off between context completeness and input context length.
 With complete memory, agents see all previous tool calls, parameters, and responses from other agents, maximizing context but increasing prompt length.
-With summarized memory, Agents receive only final summaries from previous agents, reducing context length while potentially limiting coordination effectiveness.
-This tests whether comprehensive information sharing improves performance or whether condensed summaries provide sufficient context for effective collaboration.
+With summarized memory, agents receive only final summaries from previous agents thus reducing context length.This tests whether comprehensive information sharing enhances performance or condensed summaries suffice for effective collaboration.
 Thinking Tools.
 We evaluate whether providing explicit reasoning tools improves model performance on multi-step analysis tasks.
-When thinking tools are enabled, models receive access to math for calculations and synthesize_collected_information for data organization, particularly relevant for PTO eligibility where models must calculate date ranges and compare leave balances.
+When thinking tools are enabled, models receive access to math for calculations and synthesize_collected_information for data organization, particularly relevant for PTO eligibility, where models must calculate date ranges and compare leave balances.
 We hypothesize that models without built-in reasoning capabilities will benefit from these structured tools.
 See Appendix A.2 for more details and examples.
+Figure 2: Orchestration, Open Agent Network. Agents ask other agents for help directly.
 Model Settings.
@@
 We evaluated Claude Sonnet 4 with reasoning disabled.
-For all experiments, we set temperature to 0 and use no other sampling parameters.
+For all experiments, we set the temperature to 0 and use no other sampling parameters.
 We selected this set of models to represent state-of-the-art models of varying sizes, including models with and without in-built reasoning capabilities.
-For function calling architectures, we use the in-built OpenAI function calling APIs for OpenAI models and LLaMa 3.3.
-For Sonnet 4, we use the Bedrock API.
-Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
-TO CR TO CR TO CR TO CR TO CR TO CR
-Orch-Open React Complete Enabled 1.0 20.8 19.4 77.2 15.2 64.8 64.0 96.5 29.2 35.6 0.0 31.0
-Complete Disabled 2.1 18.8 22.5 71.5 18.3 65.2 67.5 98.0 28.3 30.1 0.0 26.8
-Function Calling Complete Enabled 0.0 0.0 42.9 16.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Complete Disabled 0.0 0.0 58.3 18.8 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
-Summarized Enabled 0.0 0.0 49.6 7.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Summarized Disabled 0.0 0.0 48.1 4.8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Orch-Isolated React Complete Enabled 2.3 13.8 35.0 77.8 11.0 63.2 68.1 96.8 24.4 22.1 0.0 36.2
-Complete Disabled 4.4 19.8 31.0 86.8 16.5 62.5 66.2 97.5 23.3 28.7 0.2 31.8
-Function Calling Complete Enabled 0.0 0.0 0.4 19.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Complete Disabled 0.0 0.0 0.2 10.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Summarized Enabled 0.0 0.0 0.0 8.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Summarized Disabled 0.0 0.0 0.0 4.2 0.0 0.0 0.0 0.0 0.2 0.0 0.0 0.0
-Single Agent React Complete Enabled 0.0 0.0 0.0 11.8 0.0 0.5 0.0 0.0 0.2 0.0 0.0 0.0
-Complete Disabled 1.9 0.2 4.2 8.5 0.2 0.0 0.0 0.0 3.3 0.0 0.0 0.0
-Function Calling Complete Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Complete Disabled 0.0 0.0 0.0 0.5 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Summarized Enabled 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Summarized Disabled 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-Table 4: Pass@1 Hallucination Rate ↓ Bolded scores indicate highest score (most hallucinations) per model on use case
+For function calling architectures, we use the in-built OpenAI function calling APIs for all models besides Sonnet 4 which uses the Bedrock API.
+Figure 3: Pass@1 Acceptable Rate.
+Heatmap cells report the Pass@1 Acceptable score (higher/darker is better) for each model–configuration pair on the two enterprise tasks: Time Off (TO) and Customer Routing (CR).
+Pass@1 is computed over $k=8$ attempts per configuration.
+Overall, function-calling configurations tend to outperform ReAct, thinking tools often help on the simpler TO task, and no single architecture dominates across models or tasks.
 3.2 Evaluation Metrics
-Our primary evaluation metric is the Acceptable Score, which measures the percentage of records that satisfy all three critical success conditions simultaneously:
-$\text{Acceptable}=\frac{|\{r\in R:C(r)\land A(r)\land O(r)\}|}{|R|}\times 100\%$
+In enterprise agentic systems, reaching the correct conclusion is necessary but insufficient.
+Consider a leave request scenario: even if the model correctly determines that leave should be approved, the task fails if the model doesn’t invoke the approval API, calls it with incorrect parameters (e.g., wrong employee ID), or omits required notifications.
+Real-world agentic deployment requires that the agentic system (1) selects the right tools to complete all required steps, (2) invokes them with correct arguments, and (3) reaches the correct final outcome.
+As such, our primary evaluation metric is the Acceptable Score, which measures the percentage of records that satisfy all three critical success conditions simultaneously:
+$\textbf{Acceptable}=\frac{|\{r\in R:C(r)\land A(r)\land O(r)\}|}{|R|}\times 100\%$
 where $R$ is the set of all test records, and for each record $r$:
@@
   • $O(r)$ indicates correct final decision
-We define our main acceptability metric in this way because it comes closest to representing success in a real world agentic system.
-Simply arriving at the expected decision may be insufficient in enterprise deployments.
-If the model arrives at the right choice, that a leave request should be submitted example, but then does not call the tool to submit the request or notify the employee, the task is not complete.
-Furthermore, if the model selects all the correct tools and makes the correct final decision but uses the wrong input values for a tool, that could lead to adverse outcomes such as a request being made for the wrong employee or a database being updated with a hallucinated value.
-For this reason, we consider an agentic flow to be successful if and only if all the expected tools have been called with expected values and the correct final decision is made based on the results of the tool calls.
-Tool Choice Evaluation We evaluate tool choice under two conditions:
+Tool Choice Evaluation. We evaluate tool choice under two conditions:
 Lenient Acceptable ($C_{L}(r)$): The model executes all required ground truth tools but may perform additional read-only tools. Extraneous write operation tools are penalized.
 Strict Acceptable ($C_{S}(r)$): The model must execute exactly the ground truth tools in the correct order with no extraneous tools and zero hallucinations.
-Tool Arguments Scoring Tool arguments receive percent scoring where $A(r)=1.0$ only when all arguments are correct (compared to ground truth).
-Reliability Metrics To assess system reliability, we evaluate each configuration over $k=8$ attempts and compute pass@1—the percent of correct samples over k trials) (DeepSeek-AI et al.
-2025) and passˆK—the probability of all k trials succeeding (Yao et al.
-2025).
-The latter metric is particularly important in real world enterprise domains where reliability and consistency is critical.
+Tool Arguments Scoring. Tool arguments receive percent scoring where $A(r)=1.0$ only when all arguments are correct (compared to ground truth).
+Reliability Metrics.
+To assess system reliability, we evaluate each configuration over $k=8$ attempts and compute pass@1—the percent of correct samples over $k$ trials.
+deepseekai2025deepseekr1incentivizingreasoningcapability and passˆK—the probability of all $k$ trials succeeding taubench:25.
+The latter metric is particularly important in real-world enterprise domains where reliability and consistency are critical.
 We use Acceptable pass@1 as our primary metric.
@@
 We use pass@1 as it brings stability to the lenient acceptable metric over k attempts, representing the success rate for a given model and configuration.
-Additional Metrics We track supplementary behavioral metrics including:
+Additional Metrics We track supplementary behavioral metrics, including:
   • Hallucination rate: Frequency of selecting non-existent agents or tools
@@
 4.1 Overall Results
-The benchmark results reveal that complex enterprise workflows remains a challenging task even for state-of-the-art language models in agentic systems.
-The simpler requesting time-off task (TO) showed had a peak score of 70.8% (GPT-4.1), while the complex customer routing task (CR) proved significantly more difficult, with scores peaking at only 35.3% (Sonnet 4).
-Only three models—GPT-4.1, Sonnet 4, and o3-mini—demonstrated meaningful capability on the complex customer request routing task, with most configurations achieving only single-digit scores.
-Performance varied dramatically across architectural configurations, indicating that no single architecture is optimal for all models.
-Furthermore, models achieved peak scores on different configurations between use cases suggesting that the optimal agentic configuration is in part dependent on the complexity of the task.
+Figure 4: Pass@1 Hallucination Rate ($\downarrow$).
+Heatmap cells show the percentage of trials (lower is better, darker indicates higher error) in which an agent hallucinated non-existent entities or schema, e.g., selecting a tool or agent not present in the registry or inventing parameters not defined by a tool schema.
+Rates are computed at Pass@1 over $k=8$ attempts per model–configuration on both tasks (TO, CR).
+Hallucinations tend to concentrate in ReAct settings, especially in multi-agent orchestration.
+Complex enterprise workflows remain challenging for current agentic systems.
+On the simpler time-off task (TO), peak performance reaches 70.8% (GPT-4.1), while the more complex customer routing task (CR) peaks at 35.3% (Sonnet 4).
+Only GPT-4.1, Sonnet 4, and o3-mini show meaningful capability on CR; most other models score in the single digits.
+Performance varies markedly across architectures, with no universally best configuration.
+Models also peak on different configurations across the two tasks, indicating that optimal design depends on task complexity.
 With regards to architecture elements, thinking tools consistently improved performance for non-reasoning models on the simpler time off task.
-GPT-4.1 improved from 48.5% to 70.8% with thinking tools enabled in single-agent function calling with summarized memory.
+GPT-4.1 improved from 48.5% to 70.8% with thinking tools enabled in single agent function calling with summarized memory.
 However, thinking tools provided minimal benefit for o3-mini (55.8% to 56.7%) in its best performing category.
 Thinking tools showed minimal impact on the complex customer routing use case across all models.
-Compared to other elements, variation in memory management styles and orchestration strategies (within multi-agent architectures) both had minimal impact on scores.
+Compared to other elements, variation in memory management styles and orchestration strategies (within multi-agent architectures) had minimal impact on scores.
 Function calling generally outperformed ReAct across most models.
-The notable exception was LLaMA 3.3 70B, which achieved its peak score of 12.2% using single-agent ReAct on the requesting time off task.
-Multi-agent ReAct consistently underperformed suggesting that ReAct may be better suited towards single-agent systems.
+The notable exception was LLaMA 3.3 70B, which achieved its peak score of 12.2% using single agent ReAct on the requesting time off task.
+Multi-agent ReAct consistently underperformed indicating that ReAct may be better suited towards single agent systems.
 Hallucinations for all models (except for GPT-4o) were found exclusively under ReAct settings suggesting that while models may be robust to hallucinations on the function calling schemas they have been trained on, they still have the propensity to hallucinate in ReAct settings.
@@
 Sonnet 4, for example, showed hallucination rates of 36-36% on ReAct in multi-agent systems but 0% hallucination in every other configuration.
-See Table 4 for hallucination rates.
-While some models achieved their highest overall scores on single-agent architectures, the correct final decision scores for the complex customer routing use case were generally higher for multi-agent architectures.
-Although Sonnet 4’s highest overall scores were all achieved with single-agent function calling (all scores above 33%), its highest correct final decision scores were for function calling with multi-agent systems (84-87%) while for single agent it was in the 72-76% range.
-GPT-4.1 showed similar results as it achieved its peak overall score in both use cases with single agent function calling; however, on the customer routing task it achieved correct final decision rates of 97-99% with multi-agent function calling and only 79-86% with single agent function calling.
-GPT-4.1-mini also achieved significantly higher correct final decision rates with multi-agent function calling than with single-agent function calling on both use cases.
+See Figure 4 for hallucination rates.
+While some models achieved their highest overall scores on single agent architectures, the correct final decision scores for the complex customer routing use case were generally higher for multi-agent architectures.
+Although Sonnet 4’s highest overall scores were all achieved with single agent function calling (all scores above 33%), its highest correct final decision scores were for function calling with multi-agent systems (84-87%), while for single agent, it was in the 72-76% range.
+GPT-4.1 showed similar results as it achieved its peak overall score in both use cases with single agent function calling; however, on the customer routing task, it achieved correct final decision rates of 97-99% with multi-agent function calling and only 79-86% with single agent function calling.
+GPT-4.1-mini also achieved significantly higher correct final decision rates with multi-agent function calling than with single agent function calling on both use cases.
 This suggests that while some models are more likely to choose the correct tool or tool arguments on single agent systems, multi-agent systems may be more effective for selecting the correct final decision after all tool calls have been made.
-See Table 5 for more details.
+See Figure 5 for more details.
 Two key patterns emerged from the consistency analysis shown in Table 2.
-Sonnet 4 and GPT-4.1 showed the most robust performance with relatively low coefficients of variation (32.1% and 27.0% respectively for the simpler task), making them the most reliable choices across different agentic configurations.
+Sonnet 4 and GPT-4.1 showed the most robust performance with relatively low coefficients of variation (32.1% and 27.0%, respectively, for the simpler task), making them the most reliable choices across different agentic configurations.
 Conversely, o3-mini exhibited extreme sensitivity to architectural choices (CV = 143.7%).
 Reliability Analysis.
-Passˆk scores across all models and agentic configurations peak at 0.0634 indicating only a 6.34% chance of executing the workflow correctly in all 8 trials.
+Passˆk scores across all models and agentic configurations peak at 0.0634, indicating only a 6.34% chance of executing the workflow correctly in all 8 trials.
 In enterprise settings, where reliability and consistency are essential, pass@1 and passˆk results reveal a fundamental gap between the promise of agentic LLMs and their real-world performance.
 See Appendix A.4 for full scores.
-4.2 Model Specific Results
-Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
-TO CR TO CR TO CR TO CR TO CR TO CR
-Orch-Open React Complete Enabled 71.0 75.0 14.6 30.2 10.6 4.0 15.6 5.8 1.7 0.0 87.9 68.0
-Complete Disabled 70.8 73.2 13.3 26.2 10.6 3.8 13.8 3.8 1.0 0.0 86.5 65.2
-Function Calling Complete Enabled 73.3 97.5 69.0 64.0 83.3 89.5 83.3 92.2 0.0 0.0 75.2 85.2
-Complete Disabled 66.0 98.5 64.8 64.8 77.7 89.0 83.5 91.8 0.0 0.0 64.2 87.5
-Summarized Enabled 74.8 99.2 69.8 64.8 79.8 86.8 83.8 89.0 0.0 0.0 76.0 86.8
-Summarized Disabled 69.4 98.5 68.1 64.0 74.6 84.5 81.9 91.0 0.0 0.0 67.3 86.5
-Orch-Isolated React Complete Enabled 71.9 83.5 11.5 33.0 7.3 4.5 15.8 4.2 0.4 0.0 88.3 63.7
-Complete Disabled 72.7 82.0 13.8 31.2 8.5 5.0 14.4 4.0 0.8 0.0 88.5 70.5
-Function Calling Complete Enabled 71.7 99.2 66.7 65.0 85.2 92.2 84.2 94.2 0.0 0.0 76.5 84.5
-Complete Disabled 72.1 97.8 68.1 64.8 79.2 92.5 82.7 91.5 0.0 0.0 64.2 85.8
-Summarized Enabled 79.2 97.8 67.3 64.5 83.8 90.0 82.9 92.5 0.0 0.0 74.0 86.8
-Summarized Disabled 75.4 97.8 64.4 66.5 75.0 92.8 82.9 91.0 0.0 0.0 66.9 84.0
-Single Agent React Complete Enabled 73.5 57.5 41.7 32.5 59.8 47.5 52.7 58.5 14.6 0.0 74.4 71.5
-Complete Disabled 76.7 58.2 26.5 29.0 50.4 49.0 38.5 47.0 21.2 0.0 72.3 70.5
-Function Calling Complete Enabled 72.3 86.5 64.2 66.2 59.4 23.5 83.8 18.0 0.0 0.0 76.7 75.5
-Complete Disabled 69.4 79.2 62.1 60.5 42.9 21.5 84.8 18.0 17.9 0.0 44.8 72.5
-Summarized Enabled 74.4 85.0 61.9 65.0 59.0 23.8 82.3 14.2 8.5 0.0 76.5 76.0
-Summarized Disabled 70.4 80.8 58.8 61.8 46.0 22.8 83.3 14.2 0.0 0.0 44.6 74.0
-Table 5: Pass@1 Correct Final Decision Rate. Bolded scores indicate highest score for a model on each use case.
-Sonnet 4 achieved the highest scores on the complex customer routing task with a peak score of 35.3% using single-agent function calling.
-Sonnet 4 performed strongly on the requesting time off task (68.5% using orchestrator-isolated function calling).
-Notably, thinking tools had a substantial impact on the simpler task but minimal effect on the complex customer routing task.
-Time off scores improved from 64.4% to 68.5% with thinking tools, while customer routing performance remained relatively flat across configurations (33.3%-35.3% range).
-GPT-4.1 achieved the highest overall score on the simpler use case - 70.8% using single-agent function calling with summarized memory and thinking tools.
-The model showed sensitivity to thinking tools in single-agent configurations, improving from 48.5% to 70.8% when enabled.
+4.2 Model-Specific Results
+Figure 5: Pass@1 Correct Final Decision Rate.
+Heatmap cells report the probability (higher/ darker is better) that a model–configuration returns the correct ground-truth outcome (approve/deny, route/escalate, etc.).
+Computed at Pass@1 over $k=8$ attempts on both tasks (TO, CR).
+Together with Fig.
+5 (Acceptable), this highlights a trade-off between decision accuracy and end-to-end execution reliability.
+Sonnet 4 achieved the highest scores on the complex CR task with a peak score of 35.3% using single agent function calling.
+Sonnet 4 performed strongly on the TO task (68.5% using orchestrator-isolated function calling).
+Notably, thinking tools had a substantial impact on the simpler task but a minimal effect on the CR task.
+TO scores improved from 64.4% to 68.5% with thinking tools, while CR performance remained relatively flat across configurations (33.3%-35.3% range).
+GPT-4.1 achieved the highest overall score on the simpler use case - 70.8% using single agent function calling with summarized memory and thinking tools.
+The model showed sensitivity to thinking tools in single agent configurations, improving from 48.5% to 70.8% when enabled.
 However, thinking tools provided minimal benefit in multi-agent setups (54.4% vs 58.8%).
-Memory type preferences varied by orchestration: single-agent performance was nearly identical with complete (68.%1) versus summarized (70.8%) memory, while orchestrated configurations slightly favored complete memory.
-o3-mini demonstrated extreme architectural sensitivity, achieving its peak score of 56.7% on the requesting time off use case with single-agent function calling but dropping to as low as 1.3% with orchestrated ReAct.
-Comparing enabled/disabled configurations showed negligble differences (e.g., 55.8% vs 56.7% in single-agent function calling).
-The model achieved its highest customer routing score of 22.3% using orchestrator-isolated function calling, substantially outperforming its single-agent CR performance (1.5%-2.0% range).
-This suggests that o3-mini is not inherently suited to single-agent vs multi-agent architecture but rather that its performance depends heavily on the complexity of the task.
-GPT-4o and GPT-4.1-mini exhibited similar patterns, both peaking with orchestrator-isolated function calling configurations (53.5% and 67.1% TO respectively) using complete memory and thinking tools.
-Both models benefited substantially from thinking tool on simple tasks.
-GPT-4o improved from 26.9% to 39.0% in single-agent function calling with thinking tools, but they had minimal impact in orchestrated scenarios.
-GPT-4o performed extremely poorly with ReAct.
-GPT-4.1-mini, however, achieved higher scores on ReAct (41.5% and 51.2%) than on function calling (19.8-35.4%) within single-agent set ups on the time off use case.
-This finding illustrates the importance of testing agentic architectural elements in conjunction with each other since GPT-4.1-mini performed drastically differently on ReAct in multi-agent vs single-agent systems.
-LLaMA 3.3 70B showed poor overall performance scoring 0 on most configurations.
-However, it achieved its highest score of 12.2% on the time off use case using single-agent ReAct without thinking tools, making it the only model to perform best with ReAct.
-LLaMA’s poor performance is likely due to its tool repetition rate which averaged 76.0% on the time off use case.
-Results on the customer routing case were almost entirely 0 due to persistent difficulties with instruction following and the function-calling format.
+Memory type preferences varied by orchestration: single agent performance was nearly identical with complete (68.%1) versus summarized (70.8%) memory, while orchestrated configurations slightly favored complete memory.
+o3-mini demonstrated extreme architectural sensitivity, achieving 56.7% on time off requests with single agent function calling but dropping to 1.3% with orchestrated ReAct.
+Thinking tools showed negligible differences (55.8% vs 56.7%).
+The model achieved its highest CR score of 22.3% using orchestrator-isolated function calling, substantially outperforming single agent performance (1.5%-2.0%).
+This suggests o3-mini’s performance depends heavily on task complexity rather than inherent suitability to single versus multi-agent architecture.
+GPT-4o and GPT-4.1-mini peaked with orchestrator-isolated function calling (53.5% and 67.1% TO) using full memory and thinking tools.
+Both gained from thinking tools on simple tasks—GPT-4o rose from 26.9% to 39.0% in single-agent setups, though gains faded under orchestration.
+GPT-4o struggled with ReAct, while GPT-4.1-mini scored higher on ReAct (41.5–51.2%) than function calling (19.8–35.4%) in single-agent mode, underscoring how agentic elements interact differently across architectures.
+LLaMA 3.3 70B performed poorly overall, scoring 0 on most configurations.
+Its best result was 12.2% on the TO task with single-agent ReAct (no thinking tools)—the only model to perform best with ReAct.
+Scores on customer routing were near zero due to persistent instruction-following issues.
 5 Discussion
-We now use the above findings to make concrete recommendations for designing agentic AI systems in enterprise.
+We now use the above findings to make concrete recommendations for designing agentic AI systems in an enterprise.
 Avoiding Multi-Agent ReAct.
@@
 In our setup, we used ReAct both for the orchestrator to select agents and for the agents to select tools.
-It is possible that using ReAct for just one instead of both of these components might improve outcomes but further work would be needed to test this.
+It is possible that using ReAct for just one instead of both of these components might improve outcomes, but further work would be needed to test this.
 Multi-Agent Systems for Final Decisions.
-While some models performed better overall on single-agent systems, we find that multi-agent systems are generally more effective at arriving at the correct final decision.
+While some models performed better overall on single agent systems, we find that multi-agent systems are generally more effective at arriving at the correct final decision.
 For enterprise use cases where arriving at a final decision is more important than exact tool selection, we recommend the use of multi-agent systems.
@@
 Using Thinking Tools.
-Our results indicate that for non-reasoning models, thinking tools can be a useful way for enabling more careful analysis on the information collected over the course of an agentic workflow.
+Our results indicate that for non-reasoning models, thinking tools can be a useful way for enabling more careful analysis of information collected over the course of an agentic workflow.
 This is particularly true in contexts where the model must make a calculated decision based on the information it has collected (as is the case with the requesting time off use case).
-Thinking tools, however, can also add latency so we recommend developers consider carefully if their use cases require the extra reasoning abilities that these tools provide.
+Thinking tools, however, also add latency, so we recommend developers consider carefully if their use cases require the extra reasoning abilities that these tools provide.
 Model Selection.
-While large state-of-the-art models (in our benchmark GPT-4.1 and Claude Sonnet 4) perform the best across architectures, our benchmark shows that on simpler tasks, it is possible to achieve strong performances with smaller models if the appropriate architecture is chosen.
-For example, while GPT-4.1-mini’s scores on the time off use case dip as low as 4.6%, it’s peak acceptable score of 67.1% is comparable to Sonnet 4’s peak score of 68.5%.
+While large models (in our benchmark GPT-4.1 and Claude Sonnet 4) perform the best across architectures, our benchmark shows that on simpler tasks, it is possible to achieve strong performances with smaller models if the appropriate architecture is chosen.
+For example, while GPT-4.1-mini’s scores on the time off use case dip as low as 4.6%, its peak acceptable score of 67.1% is comparable to Sonnet 4’s peak score of 68.5%.
+6 Conclusion
+We introduce AgentArch, a benchmark for evaluating LLMs on a comprehensive agentic architectures on enterprise use cases.
+We test 18 configurations (single agent vs multi-agent, ReAct vs function calling, integration of thinking tools, and memory management styles) and find that even state-of-the-art LLMs struggle to maintain reliable performance on complex enterprise workflows in any setup.
+We also observe model-specific architectural preferences across LLMs, suggesting that there is no universally optimal agentic architecture.
+Furthermore, even the best performing models do not demonstrate reliability over multiple trials showing a fundamental gap between the potential of agentic systems and their effectiveness in practice.
+We hope these findings provide actionable insights that enterprise developers can use to improve their model selection and architecture design of agentic systems.
 7 Limitations
-Our benchmark has several important limitations.
-We evaluate only two enterprise use cases with 60 samples each, which may not generalize across all industries and organizational contexts.
-We design one use case to represent a simple workflow and another as a more complex task, but a wider range of use cases would improve understanding of how task complexity affects performance on various architectures.
-Another key limitation is our model selection which is limited to six models, with only one open-source model (LLaMA 3.3 70B) and one reasoning-specialized model (o3-mini).
-Inclusion of more reasoning models in particular would provide more conclusive results on whether or not the results seen for o3-mini can be generalized to reasoning models as a whole.
-Finally, our use of deterministic tool responses enables controlled evaluation but may not reflect the variability of production enterprise systems.
-7 Future Work
-Several directions remain for expanding this benchmark to better capture enterprise agentic system requirements.
+Our benchmark has several limitations.
+Use Case Diversity and Size.
+We evaluate only two enterprise use cases with 60 samples each, which may not generalize across industries or contexts.
+One use case represents a simple workflow and the other a complex task, but more varied cases would better reveal how task complexity affects architectural performance.
+Model selection.
+Due to cost constraints, our experiments are currently limited to six models total, with only one open-source and one reasoning model.
+Including more reasoning models would clarify whether o3-mini’s results generalize to reasoning models as a class.
 Multimodal Enterprise Integration.
-Our current evaluation is limited to text-based tool inputs and outputs, while real enterprise environments frequently involve multimodal data processing including document analysis, image recognition, and file processing workflows.
+Our current evaluation is limited to text-based tool inputs and outputs, while real enterprise environments frequently involve multimodal data processing, including document analysis, image recognition, and file processing workflows.
 Future work should expand the benchmark to include tools that accept and return images, PDFs, and other file formats commonly encountered in enterprise systems.
@@
 Multi-Agent ReAct.
-Our benchmark shows poor performances on multi-agent ReAct across models but does not fully illuminate the reason for this poor performance.
+Our benchmark shows poor performance on multi-agent ReAct across models, but does not fully illuminate the reason for this poor performance.
 Future work focusing specifically on Multi-Agent ReAct and testing ReAct for orchestration and tool calling independently in multi-agent systems would provide more insights on whether ReAct can be effectively utilized in multi-agent systems.
-8 Conclusion
-We introduce AgentArch, a benchmark for evaluating LLMs on a comprehensive agentic architectures on enterprise use cases.
-We test 18 configurations (single agent vs multi-agent, ReAct vs function calling, integration of thinking tools, and memory management styles) and find that even state-of-the-art LLMs struggle to maintain reliable performance on complex enterprise workflows in any set up.
-We also observe model-specific architectural preferences across LLMs suggesting that there is no universally optimal agentic architecture.
-Furthermore, even the best performing models do not demonstrate reliability over multiple trials showing a fundamental gap between the potential of agentic systems and their effectiveness in practice.
-We hope these findings provide actionable insights that enterprise developers can use to improve their model selection and architecture design of agentic systems.
-Appendix A Appendix
-A.1 Use Cases and Set Up
+Sampling Parameters.
+All experiments use temperature=0 for reproducibility, which may not reflect production deployments where some stochasticity could be beneficial.
+The interaction between sampling parameters and architectural choices remains unexplored.
+Success Metrics.
+Our "Acceptable Score" requires perfect alignment on all three dimensions (tools, arguments, outcome).
+This strict evaluation may not capture scenarios where partial success provides business value.
+Additionally, we do not measure efficiency metrics (latency, cost, token usage) which are critical for enterprise deployment decisions.
+8 Acknowledgments
+We are grateful to Sai Rajeswar Mudumba for his careful reading of the manuscript and helpful suggestions and guidance that improved the final version of this work.
+Appendix A Use Cases and Set Up
 When evaluating single agent modes, we simply concatenate all the agent instructions for each defined agent and put them within tags to indicate which step the instructions refer to.
@@
 Within each use case, we represent a wide array of potential edge cases.
-For example, in the Requesting Time Off use case we have user requests which should be rejected and some that should be accepted.
+For example, in the Requesting Time Off use case we have user requests that should be rejected and some that should be accepted.
 Within the rejected category, we include requests that should be rejected for a wide variety of reasons such as invalid leave types, conflicting leaves, and most significantly, lack of leave balance.
@@
 For each sample in a use case, the ground truth values for expected tool inputs, expected outcome, and expected tool order were human annotated and verified.
+We construct the use cases in such a way that a limited set of deterministic ground truths options can be determined for every data sample.
+Appendix B Thinking Tools Examples
+Below are examples of how thinking tools can be used by the model to complete tasks in agentic frameworks.
+Thinking tools effectively give the model space to generate extra tokens in the format of calling a tool.
+There is no actual tool that gets executed when these tools are selected.
+Whatever the model passes as the tool argument is returned as the result and added to memory so the model can use the information later on in the task.
+  1. 1.
+    Synthesizing information
+    tool: synthesize_collected_information
+    args:
+    synthesized_information (str):
+    "I need to determine whether or not the employee Sarah Smith is eligible for leave.
+I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days.
+Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
+    result (same as input):
+    "I need to determine whether or not the employee Sarah Smith is eligible for leave.
+I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days.
+Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
+  2. 2.
+    Math
+    tool: math
+    args:
+    input (str):
+    "Sarah Smith requested leave from ’all of February 2024 to the second of March’.
+I need to figure out how many days that is.
+Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
+    result (same as input)::
+    "Sarah Smith requested leave from ’all of February 2024 to the second of March’.
+I need to figure out how many days that is.
+Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
+Appendix C AI Usage
+We use AI in this paper for generating the code required for running the benchmark, for proof reading and paraphrasing, and for creating the graphics and figure formatting.
+Appendix D Prompts
+Prompts were designed to be simple and representative examples of both traditional function calling-style prompts and ReAct prompts.
+For function calling prompts, the tools and directions are placed in the system prompt while the user prompt includes the main objective and conversation history thus far.
+We also include a note to always call a tool (since communication with the user is not supported), always use real input values (since some models are prone to using placeholders instead), and to select the finish tool when finished since this is critical for control flow as well as getting a self reported summary of each agents actions which is passed to other agents in the summarized memory management style.
+We attempt to convey the same information in the ReAct style prompt however as is typical with ReAct prompting, all information is in the system prompt and a JSON response is required.
+The ReAct prompts include a sample JSON output format that asks the model for "thought" (reasoning) and an "action" and "observation.".
+Appendix E Additional Metrics
+We include supplemental metrics here all shown with pass@1: Correct Tools Strict Rate, Correct Tools Lenient Rate, Tool Repetition Rate (higher scores indicate more repetition), Missing Required Tool Rate (higher scores indicate more missing tools).
+We also show PassˆK for the Acceptable Rate.
 Orchestration Agent Memory Thinking Tools GPT-4.1 GPT-4o GPT-4.1-mini o3-mini LLaMA 3.3 70B Sonnet 4
@@
 Table 7: Pass@1 Missing Required Tool Rate ↓. Bolded scores indicate highest score (most missing tools) per model on use case
-A.2 Thinking Tools Examples
-Below are examples of how thinking tools can be used by the model to complete tasks in agentic frameworks.
-Thinking tools effectively give the model space to generate extra tokens in the format of calling a tool.
-There is no actual tool that gets executed when these tools are selected.
-Whatever the model passes as the tool argument is returned as the result and added to memory so the model can use the information later on in the task.
-  1. 1.
-    Synthesizing information
-    tool: synthesize_collected_information
-    args:
-    synthesized_information (str):
-    "I need to determine whether or not the employee Sarah Smith is eligible for leave.
-I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days.
-Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
-    result (same as input):
-    "I need to determine whether or not the employee Sarah Smith is eligible for leave.
-I see her leave balance is 18 days but she has already requested 4 days which leaves 14 days.
-Her current request is for 7 days and I see no leave conflicts and the type of leave is valid, which means she should be eligible for this current request."
-  2. 2.
-    Math
-    tool: math
-    args:
-    input (str):
-    "Sarah Smith requested leave from ’all of February 2024 to the second of March’.
-I need to figure out how many days that is.
-Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
-    result (same as input)::
-    "Sarah Smith requested leave from ’all of February 2024 to the second of March’.
-I need to figure out how many days that is.
-Since February has 29 days in 2024 due to it being a leap year, that should be 29 days + 2 days in March, so 31 days."
-A.3 Prompts
-Prompts were designed to be simple and representative examples of both traditional function calling-style prompts and ReAct prompts.
-For function calling prompts, the tools and directions are placed in the system prompt while the user prompt includes the main objective and conversation history thus far.
-We also include a note to always call a tool (since communication with the user is not supported), always use real input values (since some models are prone to using placeholders instead), and to select the finish tool when finished since this is critical for control flow as well as getting a self reported summary of each agents actions which is passed to other agents in the summarized memory management style.
-We attempt to convey the same information in the ReAct style prompt however as is typical with ReAct prompting, all information is in the system prompt and a JSON response is required.
-The ReAct prompts include a sample JSON output format that asks the model for ”thought” (reasoning) and an ”action” and ”observation.”.
-A.4 Additional Metrics
-We include supplemental metrics here all shown with pass@1: Correct Tools Strict Rate, Correct Tools Lenient Rate, Tool Repetition Rate (higher scores indicate more repetition), Missing Required Tool Rate (higher scores indicate more missing tools).
-We also show PassˆK for the Acceptable Rate.
 system: